Verified Answer :
Answer: 0.54
Step-by-step explanation:
We know that there are 12 face cards out of the 52 cards of a standard deck. So the probability of drawing a face card will be:
\(\text{ P(face)}=\frac{12}{52}=\frac{3}{13}\)
Logically the probability of not drawing a face card will be:
\(\text{ P(not face)}=1-\text{ P(face)}=\frac{10}{13}\)
Since the drawn cards by audience member are put back in the deck, each draw is independent from the previous draw.
Therefore, the probability that of the 3 cards drawn, no card is is a face card is given by :
\(\text{P(not face card)}=\frac{10}{13}\times\frac{10}{13}\times\frac{10}{13}=\frac{1000}{2197}\)
Now, the probability that of the 3 cards drawn, at least 1 is a face card is given by :-
\(\text{P(atleast one face card)}=1-\text{P(not face card)}=1-\frac{1000}{2197}=\frac{1197}{2197}=0.54483\approx0.54\)
Using the binomial distribution, it is found that there is a 0.545 = 54.5% probability that of the 3 cards drawn, at least 1 is a face card.
For each card drawn, there are only two possible outcomes, either it is a face card, or it is not. The probability of a card being a face card is independent of any other card, which means that the binomial distribution is used to solve this question.
Binomial probability distribution
\(P(X = x) = C_{n,x}.p^{x}.(1-p)^{n-x}\)
\(C_{n,x} = \frac{n!}{x!(n-x)!}\)
The parameters are:
- x is the number of successes.
- n is the number of trials.
- p is the probability of a success on a single trial.
In this problem:
- 3 cards are drawn, thus \(n = 3\).
- In a deck of 52 cards, 12 are faces, thus \(p = \frac{12}{52} = 0.2308\).
The probability of at least one face is given by:
\(P(X \geq 1) = 1 - P(X = 0)\)
In which:
\(P(X = x) = C_{n,x}.p^{x}.(1-p)^{n-x}\)
\(P(X = 0) = C_{3,0}.(0.2308)^{0}.(0.7692)^{3} = 0.455\)
Then:
\(P(X \geq 1) = 1 - P(X = 0) = 1 - 0.455 = 0.545\)
0.545 = 54.5% probability that of the 3 cards drawn, at least 1 is a face card.
A similar problem is given at https://brainly.com/question/24863377